OpenStudy (anonymous):

Choose the correct x-intercept and y-intercept of the equation x + 2y = 4. (2 points) x-intercept: (0, 2), y-intercept: (4, 0) x-intercept: (2, 0), y-intercept: (4, 0) x-intercept: (0, 4), y-intercept: (2, 0) x-intercept: (4, 0), y-intercept: (0, 2)

OpenStudy (lukebluefive):

Rewrite the equation in terms of x only: \[x + 2y = 4\] \[2y = 4 - x\] \[y = -\frac{1}{2}x + 2\] This is general equation that is slightly different from slope-intercept form: \[y - a = mx - b\] Where a is the y-intercept and b is the x-intercept. Can you take it from here?

OpenStudy (lukebluefive):

Forgot to mention that m is the slope.